The function, p(d) = 1 + d/33, gives the pressure, in atmospheres (atm), at a
depth d in the sea (d is in feet). Note that p(0) = 1 atm, p(33) = 2, and so on. Find the pressure at 120 feet. The pressure at 120 feet is ___ atm.
(type an integer or a simplified fraction)

Found 2 solutions by solver91311, richard1234:
Answer by solver91311(24713)



Substitute 120 in place of d in the function and then do the arithmetic.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by richard1234(7193)

p(120) = 1 + 120/33 = 153/33 = 51/11 atmospheres.